OpenStudy (miss.rose):

@Ghostgate @zoeeh127 @JoeDeWise @ShiBuringa Can you help me with 12 math questions? I will medal!

OpenStudy (anonymous):

yes

OpenStudy (shiburinga):

Will do my best :)

OpenStudy (miss.rose):

The first question is: Which word best describes the following math sentence? 5x + 4 = 24 A. Equation B. Inequality C. Absolute Value D. Expression Please help!

OpenStudy (pawanyadav):

As it contains sign of equality so its an equation.

OpenStudy (shiburinga):

Lol You can cross out C and B.

OpenStudy (shiburinga):

@Pawanyadav try to give them hints >.<

OpenStudy (pawanyadav):

Its everything

OpenStudy (shiburinga):

Now I cant ask hich you think it is... e.e

OpenStudy (shiburinga):

Next? :o

OpenStudy (miss.rose):

So its an equation? Next Question: Which lists all the integer solutions of the equation |x| = 10? A. –10 and 10 I think it is this one. But I'm open to learning. B. 10 only C. 0 and 10 D. -10 only

OpenStudy (shiburinga):

Yep :]

OpenStudy (pawanyadav):

You think absolutely right

OpenStudy (shiburinga):

Lool when it's like |x| it's always the opposite.. xP

OpenStudy (anonymous):

The first question is not b or d

OpenStudy (shiburinga):

Wait I got it backwards for #2 >.<

OpenStudy (shiburinga):

It's always non-negative.

OpenStudy (miss.rose):

Ready for the next question? Here it is: Which lists all the integer solutions of the inequality |x| < 4? A. -3, -2, -1, 0, 1, 2, and 3 This one? B. 0, 1, 2, and 3 C. 0, 1, 2, 3, and 4 D. -4, -3, -2 , -1, 0, 1, 2, 3, 4

OpenStudy (shiburinga):

It's definitley not D.. nor C.

OpenStudy (shiburinga):

Why do you think its A?

OpenStudy (miss.rose):

Because all of the numbers listed are less than 4.

OpenStudy (shiburinga):

Mhm.. but so is B. xD but because it has more numbers..? xP

OpenStudy (miss.rose):

Well, negative numbers can work in that equation.

OpenStudy (shiburinga):

Mhm. I agree with A. xP

OpenStudy (pawanyadav):

That was asking for all solutions so It's A

OpenStudy (shiburinga):

Mhm.

OpenStudy (miss.rose):

Next Question: Yesterday, Jed's bank account balance was –$26.34. He deposited some money into his account so that the balance is now $42.18. How much did he deposit into his account? A. $68. 52 B. $68.42 C. $16.84 D. $15.48

OpenStudy (shiburinga):

What do you think it is?

OpenStudy (shiburinga):

To me it's pretty obvious.. xP

OpenStudy (miss.rose):

A?

OpenStudy (shiburinga):

Mhm. xP -26.34+42.18=? xP

OpenStudy (shiburinga):

Lol next? xP

OpenStudy (miss.rose):

Next Question: What value of k makes the statement true? K – 12 = 6 + (–12) A. -18 B. -6 C. 6 This one? D. 18

OpenStudy (pawanyadav):

Correct ...

OpenStudy (pawanyadav):

Just compare LHS and RHS

OpenStudy (shiburinga):

Lol hes right and u are to

OpenStudy (shiburinga):

brb :)

OpenStudy (miss.rose):

Oh... Ok. Thanks! Next: Solve the word problem. Luke mowed a lawn today to earn some extra money. He spent $2.10 on fuel and $4.45 on lunch. He charged $13.50 to mow the lawn. What was his profit for mowing the lawn? A. $6.95 This one? B. $20.05 C. $11.15 D. $15.85

OpenStudy (shiburinga):

Mhm. xP

OpenStudy (shiburinga):

What grade is this? It's so easy.. xD

OpenStudy (miss.rose):

7th. What grade are you in? So are you ready for the next question? 1. A biologist is recording the loss of fish in a pond. He notes the number of fish, f, in the pond on June 1. On July 1 there were 63 fish in the pond, which is 52 fewer fish than were in the pond on June 1. Which equation represents this situation? A. f + 63 = -52 B. 63 - f = 52 C. f + 63 = 52 D. f - 52 = 63 This one?

OpenStudy (shiburinga):

Oh sorry was busy ^.^

OpenStudy (shiburinga):

Okay... I agree with ur answer. :)

OpenStudy (miss.rose):

Okay. I am about to submit it. I will be right back. Then I am going to come back on give you a medal and close the question. Thanks for all your help!!!!!!

OpenStudy (shiburinga):

Lol no need to medal me. :)

OpenStudy (miss.rose):

But you deserve it. You helped me a lot.

OpenStudy (shiburinga):

Loool My friend fave me alot of medals today. xD

OpenStudy (miss.rose):

Oh... Well, lucky you. Hey I am going to have some more questions do you think you can help me?

OpenStudy (ghostgate):

I'm here now, so do you still have questions?

OpenStudy (miss.rose):

No. But thanks anyways.

OpenStudy (ghostgate):

Oh, okay. Have a great day! ;)